# Violation of Quantum Bilocal Inequalities on Mutually-Commuting von Neumann Algebra Models

### 01 Apr 2026 v2

@@ 5. Maximal violation of bilocal inequalities and algebraic structures
and implies that $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=\tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$ .
[…]
The above Theorem 5 indicates that if the bilocal inequality is maximally violated, then the underlying algebra locally contains an $M_{2}(\mathbb{C})$ factor. In contrast, the follwing theorem presents a highly specialized scenario in which a precise characterization of algebras attaining maximal viloation can be obtained. Note that next we denote by $\Lambda_{1}\overline{\otimes}\Lambda_{2}$ the von Neumann algebra of the tensor product of $\Lambda_{1},\Lambda_{2}$
Theorem 6. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, if there are hyperfinite type II1 factors $\mathcal{R}_{\mathcal{A}}$ and $\mathcal{R}_{\mathcal{C}}$ such that $\mathcal{M}_{\mathcal{A}}\simeq\mathcal{M}_{\mathcal{A}}\overline{\otimes}\mathcal{R}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}\simeq\mathcal{M}_{\mathcal{C}}\overline{\otimes}\mathcal{R}_{\mathcal{C}}$, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$.
Proof. We first deal with the special case: $\mathcal{M}_{\mathcal{A}}=\mathcal{R}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}=\mathcal{R}_{\mathcal{C}}$. Let $\widetilde{\mathcal{H}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{H}_{\alpha},\Omega_{\alpha})$ be the incomplete tensor product of $\mathcal{H}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Omega_{\alpha}=\frac{1}{\sqrt{2}}(\Phi_{1}\otimes\Phi_{1}+\Phi_{2}\otimes\Phi_{2})\equiv\Omega\text{ for some basis }\{\Phi_{1},\Phi_{2}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then by the construction of Araki and Woods [79], $\mathcal{R}_{\mathcal{A}}\simeq\mathcal{R}(\mathcal{H}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Omega_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{H}}$. Similarly, let $\widetilde{\mathcal{K}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{K}_{\alpha},\Lambda_{\alpha})$ be the incomplete tensor product of $\mathcal{K}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Lambda_{\alpha}=\frac{1}{\sqrt{2}}(\Psi_{3}\otimes\Psi_{3}+\Psi_{4}\otimes\Psi_{4})\equiv\Lambda\text{ for some basis }\{\Psi_{3},\Psi_{4}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then $\mathcal{R}_{\mathcal{C}}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{K}}$. We can obtain that there is $\mathcal{R}_{\mathcal{C}}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$. Now, from Theorem 5, we know that $A_{0},\ A_{1}$ and $A_{2}=-\frac{i}{2}[A_{0},A_{1}]$ form a realization of the Pauli spin matrices on $\mathcal{H}_{\alpha}$, and the same holds for $C_{0},\ C_{1}$ and $C_{2}=-\frac{i}{2}[C_{0},C_{1}]$ in $\mathcal{K}_{\alpha}$, and $B_{0},B_{1}$ with $B_{i}^{2}=I$ such that the quantity $\mathcal{S}_{\tau}$ in Ineq. (4.1) arrives at $2\sqrt{2}$. Applying similar technologies of Eqs. (6)-(8) of Ref.[77], one can obtain that
(5.9) $\sqrt{|\langle\Omega\otimes\Lambda,[(A_{0}+A_{1})B_{0}(C_{0}+C_{1})]\Omega\otimes\Lambda\rangle|}=\sqrt{2},$
(5.10) $\sqrt{|\langle\Omega\otimes\Lambda,[(A_{0}-A_{1})B_{1}(C_{0}-C_{1})]\Omega\otimes\Lambda\rangle|}=\sqrt{2}.$
[…]
$\widetilde{T}_{\alpha}=(\widetilde{A}_{0,\alpha}+\widetilde{A}_{1,\alpha})\widetilde{B}_{0,\alpha}(\widetilde{C}_{0,\alpha}+\widetilde{C}_{1,\alpha}),$
$\widetilde{S}_{\alpha}=(\widetilde{A}_{0,\alpha}-\widetilde{A}_{1,\alpha})\widetilde{B}_{1,\alpha}(\widetilde{C}_{0,\alpha}-\widetilde{C}_{1,\alpha}).$
Furthermore, $\Xi=\bigotimes_{\alpha\in\mathbb{N}}\Xi_{\alpha},\Upsilon=\bigotimes_{\alpha\in\mathbb{N}}\Upsilon_{\alpha},$ with $\Xi_{\alpha}=\Omega_{\alpha}\otimes\Omega_{\alpha}$ and $\Upsilon_{\alpha}=\Lambda_{\alpha}\otimes\Lambda_{\alpha}$ for almost all $\alpha$. By the definition of the incomplete infinite tensor product, such vectors are total in $\widetilde{\mathcal{H}}\otimes\widetilde{\mathcal{K}}$. But one has
$\sqrt{\langle\Xi,\widetilde{T}_{\alpha}\Upsilon\rangle}=\langle\Xi,\Upsilon\rangle\cdot\sqrt{2},\sqrt{\langle\Xi,\widetilde{S}_{\alpha}\Upsilon\rangle}=\langle\Xi,\Upsilon\rangle\cdot\sqrt{2}.$
for all $\alpha$ larger than some $\alpha_{0}$ (depending on $\Xi$ and $\Upsilon$).
[…]
It is mentioned that the results of Theorem 6 can be applied to quantum field theory. In the algebraic framework of the QFT, wedge algebras (von Neumann algebras associated to specific unbounded regions in Minkowski spacetime, such as $\mathcal{O}$ associated to tangent regions, are typically type III1 factors. By the classical result of A. Connes, all injective infinite factors (except some type III0) are strongly stable (meaning they can absorb a hyperfinite II1 factor: $\mathcal{M}\simeq\mathcal{M}\overline{\otimes}\mathcal{R}_{1}$).
Corollary 7. Let $\mathcal{O}_{i}$ $(i=1,2,3)$ be wedge-shaped regions, $\mathcal{A}(\mathcal{O}_{i})$ be the wedge algebra on a separable Hilbert space $H$ with cyclic and separating vector. If $\mathcal{A}(\mathcal{O}_{1})$ and $\mathcal{A}(\mathcal{O}_{3})$ are both type III1 factors, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$ on $(\mathcal{A}(\mathcal{O}_{1})\vee\mathcal{A}(\mathcal{O}_{2})\vee\mathcal{A}(\mathcal{O}_{3}))^{\prime\prime}$.
Acknowledgement. Authors thank for your comments. This work is supported by the National Natural Science Foundation of China (Grant No. 12271394).

### 15 Apr 2026 v5

@@ 4 Maximal violation of bilocal inequalities and algebraic structures
and implies that $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=\tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$ for any $A\in\mathcal{M}_{\mathcal{A}},\ C\in\mathcal{M}_{\mathcal{C}}$.
It is straightforward to prove the converse process, as we check that Eqs. (4.1)-(4.7) hold if $\tau(A_{i}^{2}A)=\tau(A),\ \tau(B_{i}^{2}B)=\tau(B),\ \tau(C_{i}^{2}C)=\tau(C)$, and $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=0,\ \tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$. We complete the proof. $\Box$
To further elucidate the algebraic relations presented in Theorem 4.1, we provide the following corollary.
Corollary 4.2.
For the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, the bilocal inequality can be maximally violated if and only if $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ contain subalgebras isomorphic to $M_{2}(\mathbb{C})$ and there exists a faithful state $\tau\in\mathcal{M}_{\mathcal{ABC}}^{*}$ that satisfies the independent condition (*): $\tau(AC)=\tau(A)\tau(C)$ for all $A\in\mathcal{M}_{\mathcal{A}},\ C\in\mathcal{M}_{\mathcal{C}}$.
Proof Note that for any von Neumann algebra, there always exists a faithful state $\tau\in\mathcal{M}_{\mathcal{ABC}}^{*}$.
$(\Leftarrow)$ Suppose $\mathcal{M}_{\mathcal{A}}$ (resp. $\mathcal{M}_{\mathcal{C}}$) contains a subalgebra $\mathcal{M}_{\mathcal{A}}^{sub}\simeq M_{2}(\mathbb{C})$ (resp. $\mathcal{M}_{\mathcal{C}}^{sub}\simeq M_{2}(\mathbb{C})$) and $\tau$ satisfies ()(*).
Then there exist operators $A_{0},A_{1},$ and $A_{2}:=-\frac{{\rm{i}}}{2}[A_{0},A_{1}]$ in $\mathcal{M}_{\mathcal{A}}^{sub}$ such that they anticommute and $A_{i}^{2}=I$ for $i\in\{0,1,2\}$, where ${\rm{i}}^{2}=-1$. Consequently, we obtain
$\tau(A_{i}^{2}A)=\tau(A),\ \tau[(A_{0}A_{1}+A_{1}A_{0})A]=0$
for any $A\in\mathcal{M}_{\mathcal{A}}$. Similarly for $C\in\mathcal{M}_{\mathcal{C}}$ with $C_{0},C_{1},C_{2}:=-\frac{{\rm{i}}}{2}[C_{0},C_{1}]$.
Setting $B_{0}=B_{1}=I$ gives $\tau(B_{i}^{2}B)=\tau(B)$ for all $B\in\mathcal{M}_{\mathcal{B}}$. By Theorem 4.1, these operators yield the maximal violation $2\sqrt{2}$ of the quantity $\mathcal{S}_{\tau}$ in Eq. (3.1).
$(\Rightarrow)$ Conversely, assume $\mathcal{S}_{\tau}$ in Eq. (3.1) attains the maximal violation $2\sqrt{2}$.
Then for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$ and $i\in\{0,1\}$, we have $\tau(A_{i}^{2}A)=\tau(A),\tau(B_{i}^{2}B)=\tau(B),\tau(C_{i}^{2})=\tau(C)$, and $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=0,\ \tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$.
Taking $A=A_{0}A_{1}+A_{1}A_{0}$ and using the faithfulness of state $\tau$ together with $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=0$, one gets
$A_{0}A_{1}+A_{1}A_{0}=0,$
i.e., $A_{0}A_{1}=-A_{1}A_{0}$. The algebra generated by $A_{0},A_{1}$ is
$\mathfrak{A}(A_{0},A_{1}):=\{\sum_{k}\alpha_{k}A_{0}^{m}A_{1}^{n}|\alpha_{k}\in\mathbb{C},m,n\in\mathbb{N}\}.$
From $\tau(A_{i}^{2}A)=\tau(A)$ and $I\leq A_{i}\leq I\ (i\in\{0,1\})$, setting $A=I$ gives $\tau(I-A_{i}^{2})=0$. Faithfulness of state $\tau$ then implies $A_{i}^{2}=I$. The same reasoning yields $B_{i}^{2}=I\ \textmd{and}\ C_{i}^{2}=I$.
Because $A_{i}^{2}=I$, one gets $\sum_{k}\alpha_{k}A_{0}^{m}A_{1}^{n}=\alpha_{0}I+\alpha_{1}A_{0}+\alpha_{2}A_{1}+\alpha_{3}A_{0}A_{1}$. Hence,
$\mathfrak{A}(A_{0},A_{1})={\rm span}\{I,A_{0},A_{1},-\frac{{\rm{i}}}{2}[A_{0},A_{1}]\}\simeq M_{2}(\mathbb{C}).$
Analogously, $\mathfrak{A}(C_{0},C_{1})\simeq M_{2}(\mathbb{C})$. So this proof is completed. $\Box$
The above results show a close connection between the maximal violation of Bell-type inequalities and algebraic structures. This connection can be more clearly demonstrated when all algebras are factors.
Corollary 4.3.
For the ternary tensor product algebra models of entanglement swapping networks, i.e., $\mathcal{M}_{\mathcal{ABC}}=\mathcal{M}_{\mathcal{A}}\otimes\mathcal{M}_{\mathcal{B}}\otimes\mathcal{M}_{\mathcal{C}}$, assume that $\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}}$ are factors. Then the bilocal inequality cannot be maximally violated if and only if either $\mathcal{M}_{\mathcal{A}}$ or $\mathcal{M}_{\mathcal{C}}$ is type I2k+1 for some finite positive integer $k$.
5 Discussion and Conclusions
We investigate Bell type inequality for entanglement swapping network in von Neumann algebraic framework, extending the paradigm that Bell violation from observable algebra structure (notably type III factors). We identify algebraic constraints governing inequality violation, linking network nonlocality to the noncommutative structure of the underlying algebras, and further show that maximal violation conditions can reverse-engineer the type classication of von Neumann algebras.
This work represents merely the beginning of a much broader inquiry. Our current model focuses primarily on the simplest nontrivial network: the entanglement swapping scenario with two independent sources. The generalization of these results to arbitrary multipartite quantum networks remains an important open problem. In more complex architectures, the interplay between multiple independent sources and the commutation relations of their associated algebras is expected to reveal even richer structures of nonlocality in networks represented by mutually-commuting von Neumann algebras.
Conflict of Interest  The authors declare no conflict of interest.
Acknowledgements.
We thank the referees for their time and comments.